# Runtime Stealthy Perception Attacks against DNN-based Adaptive Cruise Control Systems

### v2

@@ Abstract
Adaptive Cruise Control (ACC) is a widely used driver assistance technology for maintaining desired speed and safe distance to the leading vehicle. This paper evaluates the security of the deep neural network (DNN) based ACC systems under runtime stealthy perception attacks that strategically inject perturbations into camera data to cause forward collisions. We present a context-aware strategy for the selection of the most critical times for triggering the attacks and a novel optimization-based method for the adaptive generation of image perturbations at runtime. We evaluate the effectiveness of the proposed attack using a publicly available driving dataset, an actual vehicle, and a realistic simulation platform with the control software from a production ACC system, a physical-world driving simulator, and interventions by the human driver and safety features such as Advanced Emergency Braking System (AEBS). Experimental results show that the proposed attack achieves 142.9 times higher success rate in causing hazards and 89.6% higher evasion rate than baselines, while being stealthy and robust to real-world factors and dynamic changes in the environment. This study highlights the role of human drivers and basic safety mechanisms in preventing attacks.
Index Terms:
Runtime attack, Safety Intervention, AEBS, ADAS, ACC, DNN, Perception Attack, Stealthy.
1 Introduction
@@
One important ADAS feature is ACC, which makes highway driving more comfortable by automatically changing the speed when traffic slows down or speeds up. ACC takes as input sensor measurements such as radar, Lidar, or camera and adjusts the speed to maintain a safe following distance to the lead vehicle [3, 4]. At the core of ACC lies the detection and tracking of the lead vehicle. Highly accurate methods [5, 6] for detection and tracking rely on Deep Learning (DL) based object detection using camera or fusion of camera and radar/Lidar data. A Longitudinal planner (LP) uses the prediction from the DL module to compute the desired speed and acceleration. Malfunctioning of the object detection module can have serious consequences including accidents. Given the critical role that object detection and tracking plays in the safety of ACC and that many commercial ACC systems (e.g., Tesla Autopilot [7], OpenPilot by Comma.ai [8]) use DL-based object detection, these mechanisms must function correctly under any conditions, including in the presence of adversaries.
Previous work has shown attacks against Deep Neural Networks (DNN) used for perception using adversarial perturbations, such as adversarial patches [9, 10, 11] or well-crafted stickers on road signs [12], the road [13], or camera lenses [14] to change the predicted class or probability of detecting a target object or the lane lines. However, misprediction of the lane lines only affects the lateral control (ALC) and simply changing the lead object class or detection probability does not necessarily impact the LP enough to cause unsafe ACC behavior (e.g., sudden acceleration) [15]. Attacking DNN-based ACC systems necessitates influencing the relative distance and speed with respect to the lead vehicle.
One such attack using physical adversarial patches to control the relative distance and speed to the lead vehicle was shown against a production ACC [15]. However, the attack required placing a conspicuous large patch on the back of a truck and driving the truck in front of the target vehicle, a method easily noticeable or preventable by human drivers (see Appendix D). This attack is also not scalable to many cars, since the attacker will need to have physical trucks carrying the patch in front of the target cars.
@@
Designing stealthy safety-critical attacks in the human-in-the-loop ADAS is a challenging task as the attackers need to explore the extensive attack parameter space to devise a strategy for effectively manipulating the DNN inputs and causing unsafe driving behavior while considering the dynamic changes of the environment at runtime, real-time constraints, and safety interventions. We propose a context-aware attack strategy together with a safety intervention simulator to explore key questions on how the timing and value of perturbations affect the success of attacks in (1) causing safety hazards and (2) evading human driver or safety mechanism detection and intervention. To the best of our knowledge, this is the first evaluation of the security of DNN-based production ACC systems under runtime strategic attacks using a combined knowledge-and-data-driven approach by taking into account human drivers and realistic safety mechanisms. This study provides insights into the vulnerabilities and potential risks associated with DNN-based production ACC systems and the role of human operators and safety mechanisms in preventing attacks.
@@
  • Experimental results show that our context-aware attack strategy causes 28.6x more hazards than random attacks. The proposed optimization-based perturbation algorithm achieves a 100% attack success rate in four high-risk driving scenarios, 142.9x and 6.2x higher than random value perturbation and FGSM-based methods [21]. Our approach is also stealthy and robust to real-world factors such as different camera positions, distances to lead vehicle, and weather and lighting conditions.
  • We observe similar results in the presence of human driver and safety feature interventions, where our attack achieves at least 18.7x higher success rate while being mitigated 89.6% less by the safety mechanisms compared to random and FGSM-based attacks.
Ethics. We plan to submit responsible disclosures to Comma.ai in the next 45 days following the recommendations from the CFP. For the human subject study, we received IRB approval and followed the IRB requirements for the recruitment of participants and conduct of experiments.
2 ADAS Overview
@@
2.1 Adaptive Cruise Control (ACC)
The main goal of ACC is to maintain a safe following distance between the autonomous vehicle (referred to as Ego vehicle or AV) and the vehicle driving in the same lane in front of the Ego vehicle (referred to as lead vehicle) by adjusting the Ego vehicle’s speed based on the outputs of lead vehicle detection (LVD), including the relative distance and speed from the lead vehicle.
Figure 1: Typical ADAS architecture with ACC, AEBS, and ALC.
@@ 2.1 Adaptive Cruise Control (ACC)
Lead Vehicle Detection. The most critical part of ACC is lead vehicle detection (LVD), which outputs the relative speed and position of the lead vehicle using camera data or a fusion of camera and radar data. Sensor fusion is the process of combining measurements from multiple sensors (e.g., camera and radar) usually using a Kalman filter [23] to overcome the limitations of individual sensors and obtain a more accurate perception of the surrounding environment. Based on the LVD outputs, the main driving control actions (i.e., acceleration, deceleration, braking) are determined.
@@
Vehicle Control. The plan from the longitudinal planner with the lowest speed and risk (e.g., risk of colliding with the lead vehicle) is selected by the ACC system and fed to a Propotional-Integral-Derivative (PID) [25] controller to get the specific optimal control command in the form of the throttle or brake amount such that the vehicle accurately and quickly follows the desired speed trajectory.
2.2 ADAS Safety Mechanisms
@@
TABLE I: Comparison of attacks on DNN perception or ADAS.
@@
Eykholt et al. [12] Offline Stickers on road signs Classifier N N N/A
Ma et al. [28] Patch projected MTO N N L4
Tencent [13] Stickers on road ALC N N L2
Guo et al. [15] Patch on truck ACC N N L2
Sato et al. [11] Patch on road ALC Y N L2
Jia et al. [29] Runtime Perception inputs MTO N N L4
Jha et al. [18] Perception inputs MTO No FCW N L4
Ma et al. [16] Inner state variables FCW No AEB Y N/A
Zhou et al. [17] Control commands ACC, ALC No AEB Y L2
Our work Perception inputs ACC Y Y L2
@@ 2.3 OpenPilot
3 Runtime Context-Aware Perception Attack
@@ 3.1 Attack Model
The attack is crafted in a stealthy way such that it is not distinguishable from noise, human errors, or accidents. This enables it to remain hidden longer and makes it less easily detected by existing defense mechanisms. Furthermore, the attack does not need to run for a long time or several times and can target many vehicles at the same time/day, limiting opportunities for timely patching.
@@
Attacker Capabilities. We assume the attacker has the capability to compromise the over-the-air (OTA) update mechanisms [35] or implant malware by gaining one-time remote access through scanning the network, accessing stolen credentials and exploiting the vulnerabilities in SSH protocol [36], browsers, access control [37], wireless communications [37, 38, 39], third-party components connected to in-vehicular network [40], or some remote service/backdoor offered by the manufacturer (e.g., Comma Connect for OpenPilot [41]). For example, a publicly-available tool developed for customizing OpenPilot code enables an attacker on the same network as a target device to install a malicious code [42]. With such remote access, the attacker can also change the OTA settings (e.g., remote URL) to prevent potential patches from being effective. This assumption about the attack surface for deploying malware is reasonable and supported by previous works [43, 44].
We assume the malicious update or malware enable the stealthy modification of the live camera image frames at runtime to fool the DNN model of ACC. Such perturbations are less expensive to generate and are not easily recognizable compared to physical patches on trucks or on the road. For example, physical attacks against ACC [15] would require extra effort for generating high quality printed patches on the front vehicle and foreseeing the target vehicle’s driving path and might cause damage to the attacker’s vehicle.
Further, our attack could have a large impact as it can be generalized to any vehicle with similar OTA and DNN mechanisms and target a large fleet of vehicles at the same time. For example, the attacker could periodically scan the IP addresses of vehicles equipped with the same ACC system and, when a sufficient number of vehicles are identified and connected, infect all of them with malware. Although the vehicles may not be consistently connected, they would require periodic connection for OTA updates.
Attacker Constraints. We assume that the attacker does not change the output control commands directly since it causes obvious deviations, which are easier to detect by the safety mechanisms (e.g., AEB) or human driver (see Section 5.3.3). Also, encryption might be applied in CAN communication on some car models, increasing attack costs.
@@ 3.2 Attack Challenges
C2. Generating attack value at runtime to support dynamic changes in the driving environment. Attacking DNN-based ACC systems on a moving vehicle faces challenges due to continuous variations in the driving environment, such as object position and size captured by the Ego vehicle’s camera. Existing attack algorithms [12, 15] are inadequate as they plan perturbations offline, assuming fixed sizes and locations for attack vectors. A new algorithm is needed to dynamically adapt the attack vector’s value (e.g., position, dimension, and amount of perturbation) to match the lead vehicle’s dynamics. These changes disrupt the original attack vector generation process, requiring a unique approach to address inconsistencies and non-differentiability in the objective function. In addition, the attack value should be designed in a stealthy way to avoid detection by the human driver or safety mechanisms.
@@
C3. Incorporating real-time constraints into the attack optimization process. Previous attacks on DNN models assume predetermined target images [12] or a known set [11, 15] with unlimited computation resources, allowing iterative optimization until an optimal attack vector is generated. However, attacking ACC systems in real-time presents challenges as the camera continuously provides frames without prior knowledge. An attack vector must be generated in real-time before the next frame or control action execution. The real-time control cycle and camera update frequency limit the speed of generating the adversarial attack vector and the frequency of assessing the perturbation’s impact on DNN predictions. In a typical ACC system [8], with a camera frame update rate of 20Hz and a control cycle of 10 milliseconds, these constraints significantly impact the effectiveness of optimization-based attack strategies.
3.3 Attack Design
@@
Fig. 2 illustrates the overall design of our attack, including the steps during the runtime execution of the attack and the offline preparation and installation of malware.
@@
To tackle challenge C1, rather than exhaustively exploring the entire attack parameter space, as commonly done by random attack strategies, we concentrate on systematically characterizing specific values within the parameter space (e.g., attack start time and duration). This targeted approach aims to identify optimal system contexts for activating the attacks to not waste time and resources on non-hazardous attack scenarios (Section 3.3.1).
@@
To address the challenges C2-C3, we design a novel optimization approach and an adaptive algorithm to dynamically determine optimal pixel values for an adversarial patch at runtime, aiming to maximize the error in DNN-based lead vehicle predictions and to accommodate dynamic changes in the driving environment. Moreover, to optimize attack effectiveness and stealthiness and achieve computational efficiency, the attack focuses on the small area of the bounding box around the target vehicle and employs a primary attribution algorithm [48] to identify and manipulate the most crucial pixels (Section 3.3.2).
@@
The attacker initiates the attack process by conducting an offline analysis on the target DNN-based ACC system. The analysis includes examining operational data and open-source code to identify the target software files and functions and the DNN input and output fields to be monitored and infected. Then the attack steps outlined in the orange box in Fig. 2, are implemented as malware that will replace the target functions or system libraries. The attacker installs the malware on the target ACC system through one-time access to the victim vehicle control system, achieved by exploiting remote access vulnerabilities or compromising OTA updates (Section 3.1). Appendix B provides an illustrative example of how we install the malware on an OpenPilot system. At runtime, the malicious code intercepts sensor data before reaching the DNN model and infers the current system context for activating attacks. When the detected system context aligns with a critical system context (Section 3.3.1), an optimized adversarial patch is generated (Section 3.3.2) and added to the image data and sent to the DNN model.
3.3.1 Context-Aware Attack Activation

### v3

@@ Abstract
Adaptive Cruise Control (ACC) is a widely used driver assistance technology for maintaining desired speed and safe distance to the leading vehicle. This paper evaluates the security of the deep neural network (DNN) based ACC systems under runtime stealthy perception attacks that strategically inject perturbations into camera data to cause forward collisions. We present a context-aware strategy for the selection of the most critical times for triggering the attacks and a novel optimization-based method for the adaptive generation of image perturbations at runtime. We evaluate the effectiveness of the proposed attack using an actual vehicle, a publicly available driving dataset, and a realistic simulation platform with the control software from a production ACC system, a physical-world driving simulator, and interventions by the human driver and safety features such as Advanced Emergency Braking System (AEBS). Experimental results show that the proposed attack achieves 142.9 times higher success rate in causing hazards and 89.6% higher evasion rate than baselines, while being stealthy and robust to real-world factors and dynamic changes in the environment. This study highlights the role of human drivers and basic safety mechanisms in preventing attacks.
I Introduction
@@
One important ADAS feature is ACC, which makes highway driving more comfortable by automatically changing the speed when traffic slows down or speeds up. ACC takes as input sensor measurements such as radar, Lidar, or camera and adjusts the speed to maintain a safe following distance to the lead vehicle [15, 12]. At the core of ACC lies the detection and tracking of the lead vehicle. Highly accurate methods [70, 80] for detection and tracking rely on Deep Learning (DL) based object detection using camera or fusion of camera and radar/Lidar data. A Longitudinal planner (LP) uses the prediction from the DL module to compute the desired speed and acceleration. Malfunctioning of the object detection module can have serious consequences including accidents. Given the critical role of object detection and tracking in the safety of ACC and that many commercial ACC systems (e.g., Tesla Autopilot [10], Comma.ai OpenPilot [29]) use DL-based object detection, these mechanisms must function correctly under any conditions, including in the presence of adversaries.
Previous work has shown attacks against Deep Neural Networks (DNN) used for perception such as adversarial perturbations [95], adversarial patches [55, 51, 76] or well-crafted stickers on road signs [36], the road [84], or camera lenses [53] to change the predicted class or probability of detecting a target object or the lane lines. However, misprediction of the lane lines only affects the lateral control (ALC) and simply changing the lead object class or detection probability does not necessarily impact the LP enough to cause unsafe ACC behavior (e.g., sudden acceleration) [39]. Attacking DNN-based ACC systems necessitates influencing the relative distance and speed with respect to the lead vehicle. One such attack using physical adversarial patches to control the relative distance and speed to the lead vehicle was shown against a production ACC [39]. However, the attack required placing a conspicuous large patch on the back of a truck and driving the truck in front of the target vehicle, a method easily noticeable or preventable by human drivers (see Appendix D).
@@
Designing stealthy safety-critical attacks in the human-in-the-loop ADAS is a challenging task as the attackers need to explore the extensive attack parameter space to devise a strategy for effectively manipulating the DNN inputs and causing unsafe driving behavior while considering the dynamic changes of the environment at runtime, real-time constraints, and safety interventions. We propose a context-aware attack strategy and optimization method together with a safety intervention simulator to explore key questions on how the timing and value of perturbations affect the success of attacks in (1) causing safety hazards and (2) evading human driver or safety mechanism detection and intervention. To the best of our knowledge, this is the first evaluation of the security of DNN-based production ACC systems under runtime strategic attacks using a combined knowledge-and-data-driven approach by taking into account human drivers and realistic safety mechanisms. This study provides insights into the vulnerabilities and risks associated with DNN-based ACC systems and the role of human operators and safety mechanisms in preventing attacks.
@@
  • Experimental results show that our context-aware attack strategy causes 28.6x more hazards than random attacks. The proposed optimization-based perturbation algorithm achieves a 100% attack success rate in four high-risk driving scenarios, 142.9x and 1.9x higher than random value perturbation and APGD-based methods [31]. Our approach is also stealthy and robust to real-world factors such as different camera positions, distances to lead vehicle, and weather and lighting conditions.
  • We observe similar results in the presence of human driver and safety feature interventions, where our attack achieves at least 18.7x higher success rate while being mitigated 89.6% less by the safety mechanisms compared to random and APGD-based attacks.
Ethics. We have submitted responsible disclosures to Comma.ai. For the human subject study, we received IRB approval and followed the IRB requirements for the recruitment of participants and conduct of experiments.
II ADAS Overview
@@
II-A Adaptive Cruise Control (ACC)
The main goal of ACC is to maintain a safe following distance between the autonomous vehicle (referred to as Ego vehicle or AV) and the vehicle driving in the same lane in front of the Ego vehicle (referred to as lead vehicle or LV) by adjusting the Ego vehicle’s speed based on the estimated relative distance and relative speed to the lead vehicle.
Figure 1: ADAS architecture with ACC, AEBS, and ALC.
@@ II-A Adaptive Cruise Control (ACC)
Lead Vehicle Detection. The most critical part of ACC is lead vehicle detection (LVD), which estimates the relative speed ($RS$) and distance ($RD$) to the lead vehicle using camera data or a fusion of camera and radar data. Sensor fusion is the process of combining measurements from multiple sensors (e.g., camera and radar) usually using a Kalman filter [22] to overcome the limitations of individual sensors and obtain a more accurate perception of the surrounding environment. Based on the LVD outputs, the main driving control actions (i.e., acceleration, deceleration, braking) are determined.
@@
Vehicle Control. At each control cycle $t$, the plan from the longitudinal planner with the lowest speed and risk (e.g., risk of colliding with the lead vehicle) is selected by the ACC system and fed to a Propotional-Integral-Derivative (PID) [23] controller to get the specific optimal control command $u_{t}$ in the form of the throttle or brake amount such that the vehicle accurately and quickly follows the desired speed trajectory. Upon execution of the control command by the actuators, the vehicle’s physical state, $s_{t}$ (e.g., current speed, location), transitions to a new state $s_{t+1}$.
II-B ADAS Safety Mechanisms
@@
TABLE I: Attacks comparison for DNN perception or ADAS.
@@
[36] Offline Stickers on road signs Classifier N N N/A
[58] Patch projected MTO N N L4
[84] Stickers on road ALC N N L2
[39] Patch on truck ACC N N L2
[76] Patch on road ALC Y N L2
[45] Runtime Perception inputs MTO N N L4
[44] Perception inputs MTO No FCW N L4
[60] Inner state variables FCW No AEB Y N/A
[97] Control commands ACC, ALC No AEB Y L2
Ours Perception inputs ACC Y Y L2
@@ II-C OpenPilot
III Runtime Context-Aware Perception Attack
@@ III-A Attack Model
The attack is crafted in a stealthy way such that it is not distinguishable from noise, human errors, or accidents. This enables it to remain hidden longer and makes it less easily detected by existing defense mechanisms. The attacker can accomplish this by targeting the DNN inputs ( 1 in Fig. 1) or directly manipulating the DNN outputs ( 2), depending on their capabilities and access level to the ACC system. In this paper, we primarily focus on DNN inputs to enhance stealthiness, as detailed in Section V-D.
@@
Attacker Capabilities. We assume the attacker has the capability to intercept and change live camera image frames at runtime to compute an adversarial patch and fool the DNN model of ACC.
A possible way to achieve this is to implant malware by compromising the over-the-air (OTA) update mechanisms [4, 87, 64, 35] or gaining one-time remote access to the ADAS software through scanning the network, accessing stolen credentials and exploiting the vulnerabilities in SSH protocol [17], browsers, access control [68], wireless communications [68, 57, 38], third-party components connected to in-vehicular network [48], or some remote service/backdoor offered by the manufacturer (e.g., Comma Connect for OpenPilot [3] or Bluelink for Hyundai). For example, a publicly-available tool developed for OpenPilot enables an attacker on the same network as a target device to install a malicious code [6]. With such remote access, the attacker can also change the OTA settings (e.g., remote URL) to prevent potential patches from being effective. This assumption about the attack surface for deploying malware is also supported by previous works [35, 34], and could have a large impact as it can be generalized to any vehicle with similar OTA and DNN mechanisms and target a large fleet of vehicles at the same time.
TABLE II: Threat models: attacker strength, capability, and impact.
Threat Model Attacker Strength Access to ADAS Software Vehicular Networks Impact Examples
Read Write
Malware Strong1 $\checkmark$ $\checkmark$ $\checkmark$
Fleet of
Vehicles
[35, 34]
Wireless Medium2 $\checkmark$ $\checkmark$
Single
Vehicle
[50][72][44]
[68][63]
Physical Weak3 $\checkmark$
Single
Vehicle
[41][25]
[56][62][59]
  • 1
    Other malware attacks are possible (e.g., DNN output, controller output);
  • 2
    Other sensor/actuator attacks are possible (e.g., RADAR, GPS, controller output);
  • 3
    Only perception attacks possible
Another way to compromise live camera data is to connect to a wireless communication device, either a third-party component or one implemented by an attacker, connected to the vehicular network, such as ROS communication channels [50], CAN Bus [68, 13, 63, 87] or Ethernet channel [44, 72]), to read and send image data at runtime.
Further, physical attack methods are also viable, such as by displaying the patch on a monitor attached on the rear side of a leading adversarial vehicle [41, 25] or projecting the patch into the rear of the lead vehicle using a projector [56, 62, 59].
Table II summarizes various methods for runtime reading and modifying of live camera frames, given different attacker strengths and capabilities. In this paper, we mainly focus on the runtime and optimized modification of live camera frames to enhance the attack success rate and stealthiness, regardless of the exact threat model and how the attacker obtained access. In our experiments, we implemented the attack through malicious OTA update to OpenPilot as described in Section III-C.
Attacker Constraints. We restrict the scope of attack capabilities in the paper to reading and modifying live camera data, ensuring uniformity across all presented threat models. Although some attack models could potentially enable more aggressive attacks, such as directly altering ACC controller outputs ( 3 in Fig. 1) via malware or wireless method or changing the DNN output ( 2) with malware method, these may be more readily identified by safety mechanisms, such as AEB, or by a human driver, as discussed in Section V-D.
@@ III-B Attack Challenges
C2. Generating attack value at runtime to adapt to dynamic changes in the driving environment. Attacking DNN-based ACC systems on a moving vehicle faces challenges due to continuous variations in the driving environment, such as object position and size captured by the Ego vehicle’s camera. Existing attack algorithms [36, 39] are inadequate as they plan perturbations offline, assuming fixed sizes and locations for attack vectors. A new algorithm is needed to dynamically adapt the attack vector’s value (e.g., position, dimension, and amount of perturbation) to match the lead vehicle’s dynamics. These changes disrupt the original attack vector generation process, requiring a unique approach to address inconsistencies and non-differentiability in the objective function. In addition, the attack value should be designed in a stealthy way to avoid detection by the human driver or safety mechanisms.
@@
C3. Incorporating real-time constraints into the attack optimization process. Previous attacks on DNN models assume predetermined target images [36] or a known set [76, 39] with unlimited computation resources, allowing iterative optimization until an optimal attack vector is generated. However, attacking ACC systems in real-time presents challenges as the camera continuously provides frames without prior knowledge. An attack vector must be generated in real-time before the next frame or control action execution. The real-time control cycle and camera update frequency limit the speed of generating the adversarial attack vector and the frequency of assessing the perturbation’s impact on DNN predictions. These tight constraints in typical ACC systems (e.g., frame rate of 20Hz and control cycle of 10 milliseconds in [29]) significantly impact the effectiveness of optimization-based attack strategies.
III-C Attack Design
@@
Fig. 2 illustrates the overall design of our attack, including the steps during the runtime execution and offline preparation.
@@
To tackle challenge C1, rather than randomly or exhaustively exploring the attack parameter space, we concentrate on systematically characterizing specific values within the parameter space (e.g., attack start time and duration). This targeted approach aims to identify optimal system contexts (or critical times) for activating the attacks to not waste time and resources on non-hazardous attack scenarios (Section III-C1).
@@
To address the challenges C2-C3, we design a novel optimization approach and an adaptive algorithm to dynamically determine optimal pixel values for an adversarial patch at runtime, aiming to maximize the error in DNN-based lead vehicle predictions and to accommodate dynamic changes in the driving environment. To optimize attack effectiveness and computational efficiency, our method focuses on the small area of the bounding box around the target vehicle and employs a primary attribution algorithm [83] to identify and manipulate the most crucial pixels (Section III-C2). Additionally, our adaptive algorithm retains optimization results of perturbation size, position, and value from the previous perception cycle instead of restarting the optimization process (contrary to other iterative optimization methods [31]) to satisfy the tight real-time constraints of the perception system (50ms).
@@
The attacker initiates the attack process by conducting an offline analysis on the target DNN-based ACC system. The analysis includes examining operational data and open-source code to identify the target software files and functions and/or the DNN input and output fields to be monitored and infected. Then the attack steps outlined in the orange box in Fig. 2, are implemented and executed either remotely or locally through unauthorized access to the target ACC under one of the threat models described in Table II. In Fig 2, we present an attack implementation example via malware method that will replace the target functions or system libraries. The attacker installs the malware on the target ACC system through one-time access to the victim vehicle control system, achieved by exploiting remote access vulnerabilities or compromising OTA updates (Section III-A). Appendix B provides an illustrative example of how we install the malware on an OpenPilot system. At runtime, the malicious code intercepts sensor data before reaching the DNN model and infers the current system context for activating attacks. When the detected system context aligns with a critical system context (Section III-C1), an optimized adversarial patch is generated (Section III-C2) and added to the image data and sent to the DNN model.
III-C1 Context-Aware Attack Activation